14. Halla la solución de la ecuación cuadrática: -3- y2 = 24 {+373) {+973) ©{+313) @{+9iV3} -27
Answer #1

tout wongb -3-9=24 =) - > = 24+3 ジーゾ=27 = y = -27 y=I3ir3 @is Cobreet am sinero.